# Mugi: Value Level Parallelism For Efficient LLMs

### v1

@@ Abstract.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Architectural Support for Programming Languages and Operating Systems; March 22–26, 2026; Pittsburgh, USAisbn: 978-X-XXXX-XXXX-X/XX/$X$
@@ 1. Introduction
First, prior VLP architectures do not support LLM nonlinear operations, such as such as SiLU (Elfwing et al., 2018), Swish (Ramachandran et al., 2017), GELU (Hendrycks and Gimpel, 2016), and softmax. These operations are far more complicated than the ReLU predecessor (Fukushima, 1969) and account for significant runtime if not optimized (Choi et al., 2022; Rakka et al., 2025; Karami et al., 2025; Stevens et al., 2021; Wu et al., 2021), despite various software (Stuart F. Oberman, 1997; Maciej Wielgosz, 2009) and hardware (Wang et al., 2018; Nilsson et al., 2014; Wang et al., 2021; Howard et al., 2019; Wu et al., 2021) solutions being proposed.
@@
Second, prior VLP architectures misalign the trending asymmetric quantization in LLM inference, offering suboptimal efficiency. Memory-intensive LLMs outgrow the memory capacity in mobile devices easily (Dubey et al., 2024; Pi, 2024; NVIDIA, 2024a). BF16-INT4 quantization has been exploited on both weight (Frantar et al., 2022) and KV cache (Hooper et al., 2024) to combat the large memory footprint while preserving accuracy. But prior Carat only supports FP8 (Pan et al., 2024).
@@
Proposal. To overcome the challenges above, we craft Mugi, a new VLP architecture that support nonlinear approximation for the first time and asymmetric, small-batch GEMM, as well as reusing the array for both nonlinear operations and GEMM for efficient LLMs. First, Mugi orchestrates the first-to-date VLP support for nonlinear approximation. Mugi approximates critical nonlinear operations in LLMs, such as softmax, SiLU, and GELU. Mugi adopts input approximation and generates a precise output for an approximate input, in contrast to common output approximation with precise input (Wang et al., 2018; Nilsson et al., 2014; Wang et al., 2021; Howard et al., 2019; Wu et al., 2021). VLP approximation is value centric and assign greater accuracy to more important inputs.
Moreover, Mugi is optimized for asymmetrically quantized, small-batch GEMMs, that are not compatible in prior VLP designs (Pan et al., 2024). LLMs leverages weight-only quantization (WOQ) (Huang et al., 2024; Ma et al., 2024; Chai et al., 2023; Lin et al., 2024; Frantar et al., 2022; Dettmers et al., 2022) for activation-weight GEMM and KV cache quantization (KVQ) (Zhang et al., 2024; Kang et al., 2024; Hooper et al., 2024; Sheng et al., 2023; Zhao et al., 2024) for activation-activation GEMM, introducing BF16-INT4 GEMM. Mugi supports such asymmetric quantization by customizing the data format and optimized mapping. This optimization ensures high utilization for both WOQ with small-batch input and KVQ with grouped query attention (GQA)  (Ainslie et al., 2023). We additionally minimize the buffer cost in Mugi over Carat via broadcasting and output buffer leaning.
@@ 2.1. Value Level Parallelism
Value level parallelism (VLP) was first proposed for GEMM operations on large-batch, low-precision data (Pan et al., 2024), with an example for vector-scalar multiplication and vector-vector outer product given in Figure 2. (a) shows the temporal coding of a variable $i$, done by a temporal converter (TC) in green, which is essentially an equivalence logic. When input $i$, of value 3, equals the number in counting-up sequence (i.e., when the counter $c$ reaches 3 in the rectangle), the TC asserts a temporal spike in red at the 3rd cycle; otherwise, no spikes is generated, indicated by the bold segments in black. (b) and (c) depicts transforming a multiplication between $i=3$ and $w=1$ into the accumulation of $w$ over time. At cycle $i$, the accumulation outputs the correct product $iw$, equivalent to $1+1+1=3$. (d) exemplifies temporal subscription. Saving the correct product (Val) into the register in yellow is enabled by the temporal spike (Sub), essentially selecting its corresponding $iw$ product of $3\times 1=3$ in red. (e) extends to scalar-vector multiplication between a scalar $w$ and a vector $\vec{i}$. The accumulation results of $w$ are shared by all vector elements. Each vector element subscribes to its own product in parallel, based on it own input (red and blue). We call such parallel, value-dependent sharing across multiple inputs as value reuse. Value reuse and temporal subscription together formulate VLP. (f) shows that a vector-vector outer product can be obtained by organizing multiple columns of scalar-vector multiplication into a 2D array. Carat maps batched input activations to rows and weights to columns, with the number of columns matching the temporal spike latency to avoid resource overprovision and maximize the resource utilization. Since the temporal spike latency increases exponentially, $2^{n}$ cycles for an $n$-bit input, it is more beneficial to keep VLP at smaller bitwidths. Therefore, Carat opts for mapping the batch dimension to rows to achieve scalable performance with large batch sizes.
@@ 2.2.1. Software Implementation
We take nonlinear operations in LLMs such as softmax, SiLU (Elfwing et al., 2018), and GELU (Hendrycks and Gimpel, 2016) as examples, formulated in Equations 12, and 3 (Paszke et al., 2019), where erf means error function. To ensure numerical stability by avoiding overflow in exp, softmax inputs are usually subtracted by the maximum of all inputs. Without KV cache, softmax can take more than $40\%$ of the total runtime in transformer models (Choi et al., 2022; Rakka et al., 2025; Karami et al., 2025; Stevens et al., 2021). The GELU function is commonly approximated as shown in Equation 4 or Equation 5 (Paszke et al., 2019). The functions easily take tens even hundreds of cycles to finished (Stuart F. Oberman, 1997; Maciej Wielgosz, 2009).
@@ 2.2.3. Taylor Series Hardware Approximation
Another popular hardware approximation is a Taylor series (Nilsson et al., 2014; Wang et al., 2021; Wu et al., 2021). The coefficient of each Taylor term is precomputed. With Horner’s rule, the computation can be transformed in to concatenated multiply-accumulate (MAC) operations (Wu et al., 2021). This transformation allows for vectorized implementation, where coefficients can be shared by all inputs efficiently. However, Taylor approximation exhibits poor accuracy when inputs are far from the Taylor expansion point. Allowing more points introduces addition hardware overheads.
@@ 2.3.2. Weight-Only Quantization
Low-bit quantization is now the de facto technique to reduce the memory footprint (Kuzmin et al., 2022). Most prior works adopt symmetric quantization, e.g., INT8 or FP8 for both inputs and weights (NVIDIA, 2024b; Kuzmin et al., 2022; Dubey et al., 2024; DeepSeek-AI et al., 2024), which are still too memory inefficient for LLMs. Developers resort to sub-byte quantization for weights, while keeping the inputs in floating point format, e.g, BF16-INT4 weight only quantization (WOQ) (Huang et al., 2024; Ma et al., 2024; Chai et al., 2023; Lin et al., 2024; Frantar et al., 2022; Dettmers et al., 2022).
@@ 2.4. Sustainable Computing
Carbon emissions have become a growing concern in the AI world ((ITU) and (WBA), 2025) and AI inference reportedly contributes up to 90% of datacenter costs (Barr, 2019). To quantify carbon emissions, carbon modeling focuses on operational carbon for workload deployment and embodied carbon for infrastructure manufacture over the full lifetime (Wu et al., 2022; Faiz et al., 2024; Patterson et al., 2021). The equivalent emissions ($\text{CO}_{2\text{eq}}$) are formulated in Equation 6, where E, CI, CPA are short for energy and carbon intensity, carbon emitted per unit area. Embodied carbon is taking over operational carbon as the majority of contributed emissions (Wu et al., 2022; Faiz et al., 2024).
@@ 5.1. Large Language Model
The evaluated LLMs are summarized in Table 1, with LLaMA2-70B supporting GQA using a group size of 8. We implement all models using the HuggingFace Transformers Package (Wolf et al., 2019), profiling and computing loss or perplexity for each model over 100 inferences. During profiling, we extract the runtime input nonlinear tensors across all tokens and record the value and exponent distributions, which is documented in Figure 4. To obtain the perplexity and loss values shown in Figure 6, we sweep various configurations for each nonlinear implementation, and show windows containing the best performing configuration. For both figures, we show the smallest and largest model of each model family, with the only exception of Llama 2 13B replacing Llama 2 70B due to memory and runtime limitations.
@@ 5.2.2. Baseline
Nonlinear approximation hardware uses alternative vector arrays with added components to implement the Taylor series and PWL approximation methods (Nilsson et al., 2014; Wang et al., 2018). The Taylor series is implemented with Horner’s method up to 9 degrees, and requires additional registers to store coefficients. PWL implementation adopts 22 segments, requiring additional registers and comparators to store and select proper segments. Both methods are configured to achieve their best perplexity as shown in Figure 6. We also consider a vector array of MAC units to precisely compute nonlinear operations, which require 44 cycles (Stuart F. Oberman, 1997; Maciej Wielgosz, 2009).
For GEMM, we compare Mugi with Carat (Pan et al., 2024), systolic arrays, SIMD arrays, FIGNA (Jang et al., 2024), and tensor cores from Nvidia Hopper GPUs (NVIDIA, 2024a; NVIDIA Corporation, 2022), as well as a Mugi-L design. Given Carat only supports FP8 GEMM with inputs mapped across rows, we modify its accumulators at the top to BF16 and map inputs across columns, while using its FP8 data path for INT4 weights. The systolic and SIMD arrays are similar, with the systolic array needing additional control hardware and a column of output accumulators, compared to SIMD’s adder trees. Like Mugi, Carat implements output stationary, while the systolic and SIMD arrays use weight stationary configurations. The FIGNA configurations extend both systolic and SIMD arrays with the FIGNA PE, which is customized for FP-INT multiplication. Additionally, scale-up versions of both systolic and SIMD arrays with MAC and FIGNA configurations are evaluated. The tensor core has GEMM shaped as 8x16x16, and is a fully pipelined design to perform 8x16x16 MAC operations per cycle. We only compare tensor core in NoC settings. The Mugi-L uses a dedicated LUT to approximate nonlinear operations, rather than temporal coding based approximation. We ensure 8 inputs share one LUT to match the throughput of Mugi. Similar to Mugi, all design’s wSRAM and oSRAM widths are selected to load the array without introducing additional latency.
@@ 5.3. Carbon Modeling
To model both Mugi and baselines operational and embodied carbon, we follow a consistent approach by previous works (Faiz et al., 2024; Wu et al., 2022; Patterson et al., 2021). Operational carbon is computed as the product of E and CI, while embodied carbon is the product of Area and CPA, both outlined in Section 2.4. For CI, we use world carbon intensity outlined in ACT (Gupta et al., 2022a). We compute CPA with $E/mm^{2}$ detailed in Dark Silicon (Brunvand et al., 2018), and convert it to CPA with the previously stated CI.
@@ 6.1. Nonlinear Approximation
Mugi contributes these gains in performance to its ability to scale to larger array sizes, sharing the compute array with GEMM. However, other designs require standalone vector arrays for nonlinear operations, where the scale is bounded by the SRAM bandwidth. Additionally, Mugi does not have to compute costly exp (Maciej Wielgosz, 2009), thus multiplier free. These improvements allow Mugi to increase throughput and efficiency compared to vector arrays for both precise and approximation configurations.
@@ 7.1. Limitations
MoE and Multi-Modal Models. Mixture-of-Experts (MoE) extend standard attention-based LLMs with selective FFN experts, selected by a softmax-based gating network (Du et al., 2021; Lepikhin et al., 2020; Dai et al., 2024). Multi-modal models support multiple modalities beyond just text (i.e., language, vision, video, etc), by either tokenizing non-language inputs (Driess et al., 2023), or combining multiple modality-specific layers (Bai et al., 2025; Team et al., 2025). There additionally exist models that leverage MoE architecture on multiple modalities (Meta AI, 2025). All these additional operations have been supported in Mugi, and different modality has been studied in this work to prove the efficacy of VLP. We conjecture Mugi should generalize to both MoE and multi-modal variants, though we leave full validation to future work.
@@ 7.2. Related Works
Nonlinear Approximation. Prior works have explored approximating nonlinear operations. Some approaches use piece-wise linear (PWL) approximations, partitioning the function into segments and selecting coefficients based on input range, while others approximate the entire function with a single simplified equation (Howard et al., 2019; Wang et al., 2018). Other works use Taylor-series expansions, which can provide high accuracy but degrade as values drift from the expansion point (Nilsson et al., 2014; Wang et al., 2021). While all approximation techniques, including Mugi, introduce some levels of approximation error, others underperform Mugi in most models while increasing area and efficiency costs, as detailed in Figure 6 and Table 3.
GEMM Acceleration. A number of prior works target GEMM acceleration. Carat first introduced VLP, enabling low-precision, large-batch value reuse for CNNs (Pan et al., 2024). While large batch sizes compliment CNN workloads, LLMs operate with smaller batch sizes and larger matrices, making Carat unsuited for such workloads. Other accelerators exploit unary computing to reduce buffer overhead using ternary RIM arrays and unary half-adders (Guo et al., 2023), improving accumulation efficiency. However, each PE still requires multipliers and accumulators, whereas Mugi eliminates costly PE MAC units altogether via VLP. Additional works identify the ability to exploit data reuse to reduce data movement. Multi-chip module designs share inputs across weight-stationary chiplets to exploit cross-chiplet reuse (Shao et al., 2019). Matrix-scaling approaches reduce large matrices into sub-matrices and scaling vectors, reducing memory transfers sharing sub-matrix computation during rescaling (Qin et al., 2025). Lastly, another work identifies that quantization reduces the number of possible inputs, enabling computation to be shared between layers where outputs do not change (Riera et al., 2018). Like Mugi, these works similarly identify data reuse techniques but largely overlook value reuse, missing further opportunities for optimization.

### v2

@@ Abstract.
value-level parallelism, value reuse, data reuse, computation reuse, unary computing, temporal coding, quantization, general matrix multiplication, nonlinear approximation, large language model, KV cache, group query attention
journalyear: 2026copyright: ccconference: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 22–26, 2026; Pittsburgh, PA, USAbooktitle: Proceedings of the 31st ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’26), March 22–26, 2026, Pittsburgh, PA, USAdoi: 10.1145/3779212.3790189isbn: 979-8-4007-2359-9/2026/03ccs: Hardware Application-specific VLSI designsccs: Hardware Emerging architecturesccs: Hardware Arithmetic and datapath circuitsccs: Computer systems organization Parallel architectures
@@ 1. Introduction
First, prior VLP architectures do not support LLM nonlinear operations, such as such as SiLU (Elfwing et al., 2017), Swish (Ramachandran et al., 2017), GELU (Hendrycks and Gimpel, 2016), and softmax. These operations are far more complicated than the ReLU predecessor (Fukushima, 1969) and account for significant runtime if not optimized (Choi et al., 2022; Rakka et al., 2025; Karami et al., 2025; Stevens et al., 2021; Wu et al., 2021a), despite various software (Oberman and Flynn, 1997; Wielgosz and Jamro, 2009) and hardware (Wang et al., 2018; Nilsson et al., 2014; Wang et al., 2021; Howard et al., 2019; Wu et al., 2021a) solutions being proposed.
@@
Second, prior VLP architectures misalign the trending asymmetric quantization in LLM inference, offering suboptimal efficiency. Memory-intensive LLMs outgrow the memory capacity in mobile devices easily (Dubey et al., 2024; Pi, 2024; NVIDIA, 2024a). BF16-INT4 quantization has been exploited on both weight (Frantar et al., 2023) and KV cache (Hooper et al., 2024) to combat the large memory footprint while preserving accuracy. But prior Carat only supports FP8 (Pan et al., 2024).
@@
Proposal. To overcome the challenges above, we craft Mugi, a new VLP architecture that support nonlinear approximation for the first time and asymmetric, small-batch GEMM, as well as reusing the array for both nonlinear operations and GEMM for efficient LLMs. First, Mugi orchestrates the first-to-date VLP support for nonlinear approximation. Mugi approximates critical nonlinear operations in LLMs, such as softmax, SiLU, and GELU. Mugi adopts input approximation and generates a precise output for an approximate input, in contrast to common output approximation with precise input (Wang et al., 2018; Nilsson et al., 2014; Wang et al., 2021; Howard et al., 2019; Wu et al., 2021a). VLP approximation is value centric and assign greater accuracy to more important inputs.
Moreover, Mugi is optimized for asymmetrically quantized, small-batch GEMMs, that are not compatible in prior VLP designs (Pan et al., 2024). LLMs leverages weight-only quantization (WOQ) (Huang et al., 2024; Ma et al., 2024; Chai et al., 2023; Lin et al., 2025; Frantar et al., 2023; Dettmers et al., 2022) for activation-weight GEMM and KV cache quantization (KVQ) (Zhang et al., 2024; Kang et al., 2024; Hooper et al., 2024; Sheng et al., 2023; Zhao et al., 2024) for activation-activation GEMM, introducing BF16-INT4 GEMM. Mugi supports such asymmetric quantization by customizing the data format and optimized mapping. This optimization ensures high utilization for both WOQ with small-batch input and KVQ with grouped query attention (GQA)  (Ainslie et al., 2023). We additionally minimize the buffer cost in Mugi over Carat via broadcasting and output buffer leaning.
@@ 2.1. Value Level Parallelism
Value level parallelism (VLP) was first proposed for GEMM operations on large-batch, low-precision data (Pan et al., 2024), with an example for vector-scalar multiplication and vector-vector outer product given in Figure 2. (a) shows the temporal coding of a variable $i$, done by a temporal converter (TC) in green, which is essentially an equivalence logic. When input $i$, of value 3, equals the number in counting-up sequence (i.e., when the counter $c$ reaches 3 in the rectangle), the TC asserts a temporal spike in red at the 3rd cycle; otherwise, no spikes is generated, indicated by the bold segments in black. (b) and (c) depicts transforming a multiplication between $i=3$ and $w=1$ into the accumulation of $w$ over time. At cycle $i$, the accumulation outputs the correct product $iw$, equivalent to $1+1+1=3$. (d) exemplifies temporal subscription. Saving the correct product (Val) into the register in yellow is enabled by the temporal spike (Sub), essentially selecting its corresponding $iw$ product of $3\times 1=3$ in red. (e) extends to scalar-vector multiplication between a scalar $w$ and a vector $\vec{i}$. The accumulation results of $w$ are shared by all vector elements. Each vector element subscribes to its own product in parallel, based on it own input (red and blue). We call such parallel, value-dependent sharing across multiple inputs as value reuse. Value reuse and temporal subscription together formulate VLP. (f) shows that a vector-vector outer product can be obtained by organizing multiple columns of scalar-vector multiplication into a 2D array. Carat maps batched input activations to rows and weights to columns, with the number of columns matching the temporal spike latency to avoid resource overprovision and maximize the resource utilization. Since the temporal spike latency increases exponentially, $2^{n}$ cycles for an $n$-bit input, it is more beneficial to keep VLP at smaller bitwidths (Wu et al., 2020, 2021d). Therefore, Carat opts for mapping the batch dimension to rows to achieve scalable performance with large batch sizes.
@@ 2.2.1. Software Implementation
We take nonlinear operations in LLMs such as softmax, SiLU (Elfwing et al., 2017), and GELU (Hendrycks and Gimpel, 2016) as examples, formulated in Equations 12, and 3 (Paszke et al., 2019), where erf means error function. To ensure numerical stability by avoiding overflow in exp, softmax inputs are usually subtracted by the maximum of all inputs. Without KV cache, softmax can take more than $40\%$ of the total runtime in transformer models (Choi et al., 2022; Rakka et al., 2025; Karami et al., 2025; Stevens et al., 2021). The GELU function is commonly approximated as shown in Equation 4 or Equation 5 (Paszke et al., 2019). The functions easily take tens even hundreds of cycles to finished (Oberman and Flynn, 1997; Wielgosz and Jamro, 2009).
@@ 2.2.3. Taylor Series Hardware Approximation
Another popular hardware approximation is a Taylor series (Nilsson et al., 2014; Wang et al., 2021; Wu et al., 2021a; Wu and Miguel, 2021; Wu et al., 2019). The coefficient of each Taylor term is precomputed. With Horner’s rule, the computation can be transformed in to concatenated multiply-accumulate (MAC) operations (Wu et al., 2021a). This transformation allows for vectorized implementation, where coefficients can be shared by all inputs efficiently. However, Taylor approximation exhibits poor accuracy when inputs are far from the Taylor expansion point. Allowing more points introduces addition hardware overheads.
@@ 2.3.2. Weight-Only Quantization
Low-bit quantization is now the de facto technique to reduce the memory footprint (Kuzmin et al., 2022). Most prior works adopt symmetric quantization, e.g., INT8 or FP8 for both inputs and weights (NVIDIA, 2024b; Kuzmin et al., 2022; Dubey et al., 2024; DeepSeek-AI et al., 2025), which are still too memory inefficient for LLMs. Developers resort to sub-byte quantization for weights, while keeping the inputs in floating point format, e.g, BF16-INT4 weight only quantization (WOQ) (Huang et al., 2024; Ma et al., 2024; Chai et al., 2023; Lin et al., 2025; Frantar et al., 2023; Dettmers et al., 2022).
@@ 2.4. Sustainable Computing
Carbon emissions have become a growing concern in the AI world ((ITU) and (WBA), 2025) and AI inference reportedly contributes up to 90% of datacenter costs (Barr, 2019). To quantify carbon emissions, carbon modeling focuses on operational carbon for workload deployment and embodied carbon for infrastructure manufacture over the full lifetime (Wu et al., 2022a; Faiz et al., 2024; Patterson et al., 2021). The equivalent emissions ($\text{CO}_{2\text{eq}}$) are formulated in Equation 6, where E, CI, CPA are short for energy and carbon intensity, carbon emitted per unit area. Embodied carbon is taking over operational carbon as the majority of contributed emissions (Wu et al., 2022a; Faiz et al., 2024).
@@ 5.1. Large Language Model
The evaluated LLMs are summarized in Table 1, with LLaMA2-70B supporting GQA using a group size of 8. We implement all models using the HuggingFace Transformers Package (Wolf et al., 2020), profiling and computing loss or perplexity for each model over 100 inferences. During profiling, we extract the runtime input nonlinear tensors across all tokens and record the value and exponent distributions, which is documented in Figure 4. To obtain the perplexity and loss values shown in Figure 6, we sweep various configurations for each nonlinear implementation, and show windows containing the best performing configuration. For both figures, we show the smallest and largest model of each model family, with the only exception of Llama 2 13B replacing Llama 2 70B due to memory and runtime limitations.
@@ 5.2.2. Baseline
Nonlinear approximation hardware uses alternative vector arrays with added components to implement the Taylor series and PWL approximation methods (Nilsson et al., 2014; Wang et al., 2018). The Taylor series is implemented with Horner’s method up to 9 degrees, and requires additional registers to store coefficients. PWL implementation adopts 22 segments, requiring additional registers and comparators to store and select proper segments. Both methods are configured to achieve their best perplexity as shown in Figure 6. We also consider a vector array of MAC units to precisely compute nonlinear operations, which require 44 cycles (Oberman and Flynn, 1997; Wielgosz and Jamro, 2009).
For GEMM, we compare Mugi with Carat (Pan et al., 2024), systolic arrays, SIMD arrays, FIGNA (Jang et al., 2024), and tensor cores from Nvidia Hopper GPUs (NVIDIA, 2024a), as well as a Mugi-L design. Given Carat only supports FP8 GEMM with inputs mapped across rows, we modify its accumulators at the top to BF16 and map inputs across columns, while using its FP8 data path for INT4 weights. The systolic and SIMD arrays are similar, with the systolic array needing additional control hardware and a column of output accumulators, compared to SIMD’s adder trees. Like Mugi, Carat implements output stationary, while the systolic and SIMD arrays use weight stationary configurations. The FIGNA configurations extend both systolic and SIMD arrays with the FIGNA PE, which is customized for FP-INT multiplication. Additionally, scale-up versions of both systolic and SIMD arrays with MAC and FIGNA configurations are evaluated. The tensor core has GEMM shaped as 8x16x16, and is a fully pipelined design to perform 8x16x16 MAC operations per cycle. We only compare tensor core in NoC settings. The Mugi-L uses a dedicated LUT to approximate nonlinear operations, rather than temporal coding based approximation. We ensure 8 inputs share one LUT to match the throughput of Mugi. Similar to Mugi, all design’s wSRAM and oSRAM widths are selected to load the array without introducing additional latency.
@@ 5.3. Carbon Modeling
To model both Mugi and baselines operational and embodied carbon, we follow a consistent approach by previous works (Faiz et al., 2024; Wu et al., 2022a; Patterson et al., 2021). Operational carbon is computed as the product of E and CI, while embodied carbon is the product of Area and CPA, both outlined in Section 2.4. For CI, we use world carbon intensity outlined in ACT (Gupta et al., 2022a). We compute CPA with $E/mm^{2}$ detailed in Dark Silicon (Brunvand et al., 2018), and convert it to CPA with the previously stated CI.
@@ 6.1. Nonlinear Approximation
Mugi contributes these gains in performance to its ability to scale to larger array sizes, sharing the compute array with GEMM. However, other designs require standalone vector arrays for nonlinear operations, where the scale is bounded by the SRAM bandwidth. Additionally, Mugi does not have to compute costly exp (Wielgosz and Jamro, 2009), thus multiplier free. These improvements allow Mugi to increase throughput and efficiency compared to vector arrays for both precise and approximation configurations.
@@ 7.1. Limitations
MoE and Multi-Modal Models. Mixture-of-Experts (MoE) extend standard attention-based LLMs with selective FFN experts, selected by a softmax-based gating network (Du et al., 2022; Lepikhin et al., 2020; Dai et al., 2024). Multi-modal models support multiple modalities beyond just text (i.e., language, vision, video, etc), by either tokenizing non-language inputs (Driess et al., 2023), or combining multiple modality-specific layers (Bai et al., 2025; Team et al., 2025). There additionally exist models that leverage MoE architecture on multiple modalities (Meta AI, 2025). All these additional operations have been supported in Mugi, and different modality has been studied in this work to prove the efficacy of VLP. We conjecture Mugi should generalize to both MoE and multi-modal variants, though we leave full validation to future work.
@@ 7.2. Related Works
Nonlinear Approximation. Prior works have explored approximating nonlinear operations. Some approaches use piece-wise linear (PWL) approximations, partitioning the function into segments and selecting coefficients based on input range, while others approximate the entire function with a single simplified equation (Howard et al., 2019; Wang et al., 2018; Wu and Miguel, 2019; Wu et al., 2021c). Other works use Taylor-series expansions, which can provide high accuracy but degrade as values drift from the expansion point (Nilsson et al., 2014; Wang et al., 2021). While all approximation techniques, including Mugi, introduce some levels of approximation error, others underperform Mugi in most models while increasing area and efficiency costs, as detailed in Figure 6 and Table 3.
GEMM Acceleration. A number of prior works target GEMM acceleration. Carat first introduced VLP, enabling low-precision, large-batch value reuse for CNNs (Pan et al., 2024). While large batch sizes compliment CNN workloads, LLMs operate with smaller batch sizes and larger matrices, making Carat unsuited for such workloads. Other accelerators exploit unary computing (Wu et al., 2020, 2021b; Wu and Miguel, 2022; Wu et al., 2022b) to reduce buffer overhead using ternary RIM arrays and unary half-adders (Guo et al., 2023), improving accumulation efficiency. However, each PE still requires multipliers and accumulators, whereas Mugi eliminates costly PE MAC units altogether via VLP. Additional works identify the ability to exploit data reuse to reduce data movement. Multi-chip module designs share inputs across weight-stationary chiplets to exploit cross-chiplet reuse (Shao et al., 2019). Matrix-scaling approaches reduce large matrices into sub-matrices and scaling vectors, reducing memory transfers sharing sub-matrix computation during rescaling (Qin et al., 2025). Lastly, another work identifies that quantization reduces the number of possible inputs, enabling computation to be shared between layers where outputs do not change (Riera et al., 2018). Like Mugi, these works similarly identify data reuse techniques but largely overlook value reuse, missing further opportunities for optimization.